Suppose that an annuity will provide for 20 annual payments of 1260 dollars, with the first payment coming 9 years from now. If the nominal rate of interest is 8.6 percent convertible monthly, what is the present value of the annuity?
Answer by Theo(13342)

the present value of the annuity is 5816.73, rounded to the nearest penny.
the 20 annual payments have a present value in the 8th year of 11545.36.
that is then brought to the present having a present value of 5816.73.
the interest rate is 8.6% compounded monthly to be equal to an effective interest rate of (1 + .086/12) ^ 12 = 1.089472134 - 1 = .089472134 * 100 = 8.9472134%.
here's what the cash flow and present value looks like in excel.